Find  and  if for every positive integer ,

a)   Ai={0,i}           b)  Ai=[-i,i].


1
Expert's answer
2021-04-15T07:21:17-0400

a)  if for every positive integer i, "A_i=\{0,i\}" , then

"\bigcup\limits_{i=1}^{+\infty}A_i=\bigcup\limits_{i=1}^{+\infty}\{0,i\}=\{0\}\cup \bigcup\limits_{i=1}^{+\infty}\{i\}=\{0\}\cup N"


"\bigcap\limits_{i=1}^{+\infty}A_i=\bigcap\limits_{i=1}^{+\infty}\{0,i\}=\{0\}"

 b) if for every positive integer i, "A_i=[-i,i]" , then

"\bigcup\limits_{i=1}^{+\infty}A_i=\bigcup\limits_{i=1}^{+\infty}[-i,i]=(-\infty,+\infty)"

since for every "x\in R" there exists "i\in R" such that "|x|\leq i", i.g. "x\in [-i,i]"


"\bigcap\limits_{i=1}^{+\infty}A_i=\bigcap\limits_{i=1}^{+\infty}[-i,i]=A_1=[-1,1]"

since "\bigcap\limits_{i=1}^{+\infty}A_i\subset A_1" and "A_1\subset A_i" for all i, which implies that "A_1\subset\bigcap\limits_{i=1}^{+\infty}A_i"